OpenStudy (anonymous):

simplify the expression. (x over x-4+6 over x+4)Divided by 3 over 2x^2-32

OpenStudy (ash2326):

\[\frac{x}{x-4+\frac{6}{x+4}}/(\frac{3}{2x^2-32})\] Is this your question?

OpenStudy (anonymous):

Idk I just know it needs to be simplified from the way it already is and I can't figure it out for an hour now .

OpenStudy (ash2326):

How's the question written in your book?

OpenStudy (anonymous):

Mialy like that . I've tried getting hold my teacher, but she's not answering right now .